# AnchorCrafter: Animate CyberAnchors Saling Your Products via Human-Object Interacting Video Generation

### v1

@@ 4.2 HOI-Appearance Perception
where $Q=W_{Q}\cdot Z_{t}$, $K_{H}=W_{KH}\cdot f_{H}$, $K_{O}=W_{KO}\cdot f_{O}$, $V_{H}=W_{VH}\cdot f_{H}$, $V_{O}=W_{VO}\cdot f_{o}$, $Z_{t}=\{z^{1:n}_{t}\}$ is the input latent frames on diffusion timestep $t$ and $W_{Q}$, $W_{KH}$, $W_{KO}$, $W_{V}$ are learnable weights for attention modules. The output features from the $HumanCA$ and $ObjectCA$ layers are then summed to produce the final feature set.
@@ 4.3 HOI-Motion Injection
Despite the appearance of preservation, another key challenge lies in generating specified interactive motions between humans and objects. In this section, three issues are considered for injecting the HOI motions.
To condition the movement of objects in video generation, a critical challenge is how to provide an unambiguous representation of the orientation and positional trajectory in 3D space. We utilize the depth map $D$ as input for the trajectory of the object. We estimate and crop the depth map of the target object, and a lightweight nine-layer convolution network is employed to process $D$. The resulting features are element-wise added to the output of the first convolutional layer of the UNet.
Building upon the control of object trajectories, we further consider the occlusion between the object and the human hands during interactions. In addition to the commonly used human skeleton $P$ for controlling the overall human pose, we extract the 3D mesh sequences $H$ of the human hands and mask the corresponding 3D mesh where the object occludes the hands. The two conditions are concatenated and processed through a convolution network with the same structure as the object interaction but with unshared parameters. The features extracted by this network are subsequently added to the features within the U-Net.
We observe that spatial differences between the input pose sequence $P$, $H$, and $D$ and the reference human image $I_{H}$ affect the generated results. For example, when the reference $I_{H}$ is an upper-body picture, and the input poses $P$ includes the full body, the inference based on the full-body pose may lead to the speculation of unseen lower-body details, resulting in distortion of the consistency and accuracy of the generated videos. To address this problem, we estimate a similarity matrix from the first frame of $P$ to the pose of $I_{H}$, and then the similarity matrix is performed on all motion conditions including $P$, $H$, and $D$ to match the spatial position with $I_{H}$.
@@ 4.4 HOI-Region Reweighting Loss
By constructing modules for HOI, our model enables control over the generated humans, objects, and their interactions. Accordingly, based on the diffusion model’s loss function as described in Eq. 1, our training objective can be formulated with $c=\{P,H,D,I_{H},I_{O}\}$. However, we observed that standard training loss caused the model to fail to adequately learn object appearances.
@@
where $S$, $S_{obj}$, $S_{hand}$ is the area of the whole image, object, and hands, $M_{inter}$ indicates the mask of the interaction region consists of object and hands, and $\eta$ is a hyperparameter. Using the inverse of the image area occupied by the interaction region as a weight, the model can assign higher training importance to the interaction area. The final loss is:
@@
After applying the final loss, pose-driven capability across diverse human appearances is preserved and objects are learned precisely.
@@ 5.1 Experimental Settings
Dataset.
We collected a human-object interaction dataset through online video collection and self-recording. The dataset consists of 44 online videos and 307 real-world videos, each approximately 25 seconds long at 30 fps. It includes interactions involving eleven individuals and 286 objects, primarily common objects such as toys, cups, and boxes. For each object, we captured reference images with a frontal view, 45° left rotation, and 45° right rotation. For each video, we utilize DWPose [38] for human skeletal pose sequences, HaMeR [23] for hand 3D mesh sequences, SAM-Track [18, 7] for object motion mask and ViTA [33] for the depth map.
Implementation Details.
Pre-trained weight of MimicMotion is used for basic pose-driven human animation. We freeze VAE, CLIP, and DINO, training all other parameters for 40K steps on four NVIDIA A100 GPUs (40GB). Due to memory limitations, our training frame number is set to 6 frames. The learning rate is $10^{-5}$ with a 300 iteration warm-up. HOI-region reweighting loss uses $\eta=1$. During inference, CFG is set to 4.0, and we apply progressive latent fusion from MimicMotion [41] for long video generation. Inferring a 6-frame video takes 20 seconds and 5.13GB VRAM.
Comparison Methods.
We compare AnchorCrafter with four SOTAs: AnimateAnyone [14], MimicMotion [41], AnyV2V [19], and AnyDoor [6]. AnimateAnyone and MimicMotion are pose-guided human video generation methods. We use pre-trained weights from the open-source implementation of AnimateAnyone https://github.com/MooreThreads/Moore-AnimateAnyone.git and provide human images with objects pasted, as these methods cannot input objects. AnyV2V is a video editing method that propagates features from the first frame, for which we provide a first frame and reference video. AnyDoor is an image editing model based on object mask positions; we train it with our object dataset, use MimicMotion for reference animation, and apply AnyDoor to edit each frame individually.
Metrics.
We evaluate across multiple dimensions, using FID [10] for image quality, FID-VID [1] and FVD [31] for video quality, and VBench [15] for subject and background consistency following HOI-Swap [37]. For object generation, we introduce Object-IoU, which uses SAM-Track to compute trajectory accuracy, and Object CLIP-Score, which measures appearance consistency by comparing generated objects to reference objects. Human pose consistency is measured using Landmark Mean Distances (LMD)[16] on body and hands with OpenPose[3].
@@ 5.2 Main Results
Figure 5: Qualitative comparisons with other methods. Different colored squares highlight the different types of generation artifacts. In the results of MimicMotion and AnimateAnyone, the objects fail to maintain their appearance and cannot move in sync with the hands, while AnyV2V generates apparent artifacts in the edited videos. The combined results of AnyDoor and MimicMotion demonstrate a lack of preservation in object details.
Method FID↓ FVD↓ FID-VID↓ Obj-IoU↑ Obj-CLIP↑ LMD (Hand)↓ LMD (Body)↓ Subj-Cons↑ Back-Cons↑
AnyV2V 234.1 2873.3 53.1 0.241 0.744 94.6 30.5 68.2 77.9
MimicMotion+AnyDoor 167.8 1668.9 37.7 0.647 0.863 13.2 26.0 93.4 90.7
AnimateAnyone 172.8 2267.0 24.2 0.361 0.832 23.2 41.5 94.9 94.1
MimicMotion 138.1 1444.9 22.3 0.411 0.876 12.1 24.3 96.3 93.9
Ours 141.7 736.5 15.0 0.848 0.919 11.7 25.7 97.4 95.3
w/o Human-Object Dual Adapter 170.5 913.0 24.2 0.802 0.886 12.1 20.9 96.4 94.9
w/o Multi-View Object Feature Fusion 177.0 1371.9 52.0 0.845 0.908 12.0 24.6 97.2 95.3
w/o 3D Hand Mesh 164.0 920.1 21.3 0.847 0.907 12.0 22.3 97.3 95.0
w/o HOI-Region Reweighting Loss 164.7 807.3 69.9 0.846 0.895 11.8 22.0 97.0 94.8
@@
Table 1: Quantitative results of our method compared with SOTAs and ablation studies. Our method significantly outperforms existing approaches in terms of numerical performance for spatial movement and appearance preservation of objects, while also matching or exceeding current methods in image and video quality, as well as human pose control capability. Subject consistency(Subj-Cons) and background consistency(Back-Cons) are percentages.
Quantitative Results.
Table 1 presents our quantitative results, demonstrating that our approach achieves superior performance in terms of video quality, object perception, and hand generation. For object motion, AnchorCrafter achieves a significantly higher Object-IoU than other methods. Although AnyDoor was provided with object location masks for each frame, it struggles to maintain the shape of the objects. Other models completely fail to produce reasonable object motions. Regarding object appearance preservation, AnchorCrafter attains superior Object-CLIP scores compared to other models. It is important to note that while AnimateAnyone and MimicMotion achieved not bad scores, this was due to them treating objects as part of the human clothing or background, keeping them static in the video, rather than as independent entities.
Qualitative Results.
Our qualitative results, illustrated in Fig. 5, showcase the performance of our method. In these figures, red boxes signify inadequate object appearance, orange boxes indicate flawed hands, and green highlight unrealistic anchors. Ours achieves superior preservation of human and object appearances, while faithfully reflecting their motions. MimicMotion generates high-quality, pose-driven human images, but objects and humans exhibit coupling issues, with the model treating objects as static parts of the clothing. AnimateAnyone fails to maintain human appearance and suffers from severe frame jitters. AnyV2V is unable to propagate the first frame to subsequent frames, completely collapsing despite being provided with a perfect first frame. The combination of MimicMotion and Anydoor maintains reasonable movements of objects and humans, but the appearance preservation is inferior to AnchorCrafter. Our method, leveraging the HOI-appearance perception and HOI-motion injection, preserves both the appearance and motion control of humans and objects, generating high-quality human-object interaction videos.
@@ 5.3 Ablation Study
Validation of HOI-Appearance Perception.
(a) w/o Dual Adapter
(b) w/o Multi-View Obj.
(c) Ours
(d) w/o 3D Hand Mesh
(e) w/o Re. Loss
(f) Ours
Figure 6: Ablation studies. Our modules improve the preservation of the object and its interactions with the hands. Multi-view obj. is multi-view object feature fusion. re. loss is HOI-region reweighting loss.
We conducted detailed ablation experiments on HOI-appearance perception. First, we ablated the human-object dual adapter, as shown in Table 1. The Object-IoU metric remains superior to other models, which can be attributed to the detailed object trajectory guidance. However, the Object-CLIP score decreased, suggesting that merely injecting the latent representation of objects is insufficient for object perception, and the coupling with human subjects leads to a decline in overall video quality. Fig. 6 presents qualitative results, demonstrating that objects are significantly influenced by human subjects and backgrounds. The HOI-appearance perception effectively learns object representations and decouples human and object features. Next, we ablated the multi-view object feature fusion by removing the links that included multi-view fusion self-attention links. As shown in Fig. 6, the model exhibits poorer texture representation of objects.
Validation of HOI-Motion Injection.
We removed the hand 3D mesh information injection from the HOI-motion injection. As shown in Table 1, the absence of detailed hand guidance resulted in a decrease in LMD (hand). Fig. 6 presents qualitative results, indicating the presence of artifacts in hand generation. This may be attributed to the complexity of hand-object interactions and the inaccuracies in hand pose recognition provided by DWPose, especially when the hand is largely occluded by objects or positioned at the edge of the frame.
Validation of HOI-Region Reweighting Loss.
In conjunction with the HOI-region reweighting loss, we assigned higher weights to the hand-object regions during training to enhance the learning of objects and hands. Fig. 6 demonstrates that without region enhancement, the model struggles to adequately learn object features.

### v2

@@ 4.2.2 Human-Object Dual Adapter
where $Q=W_{Q}\cdot Z_{t}$, $K_{H}=W_{KH}\cdot f_{H}$, $K_{O}=W_{KO}\cdot f_{O}$, $V_{H}=W_{VH}\cdot f_{H}$, $V_{O}=W_{VO}\cdot f_{o}$, $Z_{t}=\{z^{1:f}_{t}\}$ is the input latent frames on diffusion timestep $t$ and $W_{Q}$, $W_{KH}$, $W_{KO}$, $W_{V}$ are learnable weights for attention modules. Obtain the final feature set by summing the output features from the $HumanCA$ and $ObjectCA$ layers. Dual adapter effectively extracts features of objects and humans while achieving the disentanglement of the two entities.
@@ 4.3 HOI-Motion Injection
We propose an HOI-motion injection module that explicitly models hands and occlusion to address the challenge of generating specified interactive motions.
A critical challenge in conditioning the movement of objects is providing an unambiguous representation of their orientation and positional trajectory in 3D space. To address this, the depth map $D$ serves as the primary input for the object’s trajectory. Estimate and crop the depth map of the target object, then process it through a lightweight nine-layer convolutional network. Subsequently, the resulting features are element-wise added to the output of the first convolutional layer of the UNet. Due to the suboptimal performance of existing pose extractors for complex hand movements, particularly under occlusion, this study further considers the occlusions occurring between hands and objects during interactions. In addition to the commonly used human skeleton $P$ for controlling the overall human pose, we extract the 3D mesh sequences $H$ of the human hands and mask the corresponding 3D mesh where the object occludes the hands. The two conditions are concatenated and processed through a convolution network with the same structure as the object interaction but with unshared parameters. The features extracted by this network are subsequently added to the features within the UNet.
We observe that spatial differences between the input pose sequence $P$, $H$, and $D$ and the reference human image $I_{H}$ affect the generated results. For instance, when reference $I_{H}$ depicts a person closer to the camera, while the input pose $P$ is farther from the camera and includes more body parts, the input pose does not correspond to the reference image, leading to distortion in the consistency and accuracy of the generated video. To address this problem, we estimate a similarity matrix from the first frame of $P$ to the pose of $I_{H}$, and then the similarity matrix is performed on all motion conditions including $P$, $H$, and $D$ to match the spatial position with $I_{H}$.
@@ 4.4 HOI-Region Reweighting Loss
By constructing modules for HOI, our model enables control over the generated humans, objects, and their interactions. Accordingly, based on the diffusion model’s loss function described in Eq. 1, our training objective can be formulated with $c=\{P,H,D,I_{H},I_{O}\}$. However, we observed that standard training loss caused the model to fail to learn object appearances adequately.
@@
where $S$, $S_{obj}$, $S_{hand}$ is the area of the whole image, object, and hands, $M_{inter}$ indicates the mask of the interaction region consisting of object and hands, and $\eta$ is a hyperparameter. Using the inverse of the image area occupied by the interaction region as a weight, the model can assign higher training importance to the interaction area. The final loss is:
@@
After applying the final loss, pose-driven capability across diverse human appearances is preserved, and objects are learned precisely.
@@ 5 Experiments
This chapter provides a comprehensive evaluation of our system. We first introduce the dataset and experimental setup (Sec. 5.1), followed by benchmarking against STOAs to assess appearance preservation, motion control, and interaction realism (Sec. 5.2). To further validate our design choices, we conduct ablation studies on HOI-appearance perception, HOI-motion injection, loss reweighting, and fine-tuning strategies (Sec. 5.3), along with a user study (Sec. 5.4). Finally, we discuss the scalability of our method, which can handle non-rigid and similarly shaped objects (Sec. 5.5).
@@ 5.1 Experimental Settings
5.1.1 Dataset
For training, due to the lack of an open-source dataset specifically suited to our task, we constructed a custom dataset comprising two types of videos: human-only videos and human-object interaction videos.
  • Human-only videos. The human-only subsets comprising a total of 9,906 videos originate from two open-source datasets: HumanVid [41] and DreamDance [43]. HumanVid includes a variety of resolutions and encompasses both internet-sourced and generated data. We filtered the HumanVid to retain only vertical videos from the internet-sourced subset featuring a single person, resulting in a collection of 5,848 videos. Meanwhile, DreamDance is a human dance dataset providing URLs for each video. Due to access restrictions, we successfully downloaded 4,058 videos.
  • Human-object interaction videos. The human-object interaction subset includes 356 real-world videos recorded using mobile devices, as well as 44 livestream recordings collected from online platforms. These videos feature 11 participants interacting with 286 distinct objects, primarily comprising common household items such as toys, cups, and boxes. Each video has a duration of approximately 20 seconds at 30 FPS, capturing natural and spontaneous human-object interactions. To ensure diverse and complete object representation, we photograph objects from three perspectives: frontal, 45° left rotation, and 45° right rotation. For a small number of objects lacking three-view coverage, duplicated images were used to approximate multi-view input.
For testing, we compiled a dataset of 80 videos, featuring interactions between eight individuals sourced from internet and five distinct objects. Each object includes approximately one minute of interaction footage for fine-tuning, along with two eight-second clips for evaluation. The individuals engage with the objects in diverse poses, ensuring a broad range of interaction scenarios. The training and testing datasets have been made public.
Pre-processing. For each video, we use DWPose [38] to extract human skeletal pose sequences, following [7]. For the videos we recorded, we additionally employ HaMeR [51] to extract hand 3D mesh sequences, SAM2 [52] to extract object motion masks, and ViTA [53] to produce depth maps. For the human-only videos, since they do not involve objects and rapid hand movements result in blurring, we use black pixels to represent objects, hand sequences, and object depth maps.
5.1.2 Implementation Details
The pre-trained weight of MimicMotion [7] is used for basic pose-driven human animation. We freeze VAE, CLIP, and DINO, training all other parameters on seven gpus with 40 GB. In the first stage, we train for 12,000 iterations at a resolution of ${512\times 768}$ with 10 frames per sample. In the second stage, we train for 4,000 iterations at ${576\times 1024}$ with six frames per sample. During these two processes, 65% of the training data is sampled from our recorded dataset, while 35% is sampled from open-source datasets. During the fine-tuning phase, we train five objects simultaneously. To enhance the generalization capability for face generation and mitigate the impact of limited human diversity, 80% of the training data is sampled from the fine-tuning dataset, while the remaining 20% comes from DreamDance. The model undergoes 3,500 iterations at a resolution of ${576\times 1024}$, requiring approximately three to four hours. The learning rate is set to $10^{-5}$ with a 300-iteration warm-up. The HOI-region reweighting loss is applied with $\eta=1$ to enhance object detail learning.
During inference, CFG is set to 4.0. And we apply the time-aware position shift fusion from Sonic [54] to generate long videos, processing ten frames per sample. Inference for a 100-frame video takes five minutes and requires 22.84 GiB of VRAM.
5.1.3 Comparison Methods
We compare AnchorCrafter with five SOTAs: MimicMotion [7], StableAnimator [55], UniAnimate-DIT [23], Vace [24], and FlexiAct [56]. MimicMotion and StableAnimator are pose-guided human video generation methods based on UNet. UniAnimate-DIT is a digital human synthesis model leveraging the DiT architecture, while Vace is an all-in-one DiT-based model designed for both video creation and editing. FlexiAct is an advanced motion transfer framework that requires fine-tuning for specific actions. Consequently, we conducted a qualitative analysis to compare its performance. As these models do not directly accept object inputs, we prepared human images with objects pasted onto them for comparisons.
5.1.4 Metrics
We evaluate our system across multiple perspectives, including video quality, object shape, and appearance understanding, ensuring a comprehensive assessment of interaction realism and detail preservation. Using VBench [57] to measure subject consistency (Subj-Cons) and background consistency (Back-Cons) as well as motion smoothness (Mot-Smth). For object generation, we introduce Object-IoU, which utilizes SAM2 to extract object motion trajectory masks and compute the intersection over union (IoU) with the ground-truth mask. Additionally, Object-CLIP evaluates appearance consistency by computing the CLIP cosine similarity between generated objects and reference objects. For human generation, we use AdaFace [58] to extract facial features and calculate the cosine similarity between the generated face and the reference face, denoted as face-cos. We measure hand consistency using Landmark Mean Distances (LMD) [8], leveraging OpenPose [26].
@@ 5.2 Main Results
Figure 6: Qualitative comparisons with other methods. Other approaches fail to preserve the visual integrity of individuals and objects or maintain reasonable interactive motion, making them incapable of completing character interaction tasks.
@@
TABLE I: The quantitative results of our method were compared with those of SOTAs and ablation studies. Our method significantly outperforms existing approaches regarding numerical performance for spatial movement and appearance preservation of objects while also matching or exceeding current methods in image and video quality and human pose control capability. Subject consistency (Subj-Cons) and background consistency (Back-Cons) are percentages.
Method Object Human Overall
Obj-IoU↑ Obj-CLIP↑ Face-Cos↑ LMD (Hand)↓ Subj-Cons↑ Back-Cons↑ Mot-Smth↑
MimicMotion 0.324 0.843 0.60 20.43 93.24 93.23 98.62
StableAnimator 0.341 0.846 0.55 19.14 92.95 91.96 98.38
UniAnimate-DiT - 0.772 0.73 - 93.73 91.73 98.88
VACE 0.411 0.857 0.40 22.78 91.50 91.44 98.55
Ours 0.906 0.921 0.70 12.19 95.43 94.96 98.91
w/o Human-Object Dual Adapter 0.908 0.892 0.62 12.33 95.30 94.75 98.91
w/o Multi-View Self-Attention 0.894 0.904 0.69 12.41 95.37 94.74 98.87
w/o HOI-Motion Injection 0.903 0.918 0.68 13.19 95.88 95.13 98.92
w/o HOI-Region Reweighting Loss 0.895 0.904 0.65 12.46 95.40 94.91 98.94
w/o Fine-Tuning 0.912 0.847 0.71 12.83 94.75 94.32 98.85
5.2.1 Quantitative Results
Table I presents the quantitative evaluation of our approach. Due to the complexity of UniAnimate-DiT’s alignment strategy, ground truth for pose is unavailable. Consequently, we did not assess the accuracy of hand and object positioning. The results highlight the superior performance of our method in object appearance reproduction, object motion, and hand generation.
In terms of object motion, we achieves a significantly higher Object-IoU compared to competing methods, while other models fail to generate plausible object trajectories. Regarding object appearance preservation, we attains the highest Object-CLIP score among all evaluated approaches. Notably, some models attained relatively high scores; however, this was largely due to their tendency to treat objects as extensions of human clothing or background, causing them to remain static rather than as distinct, independent entities.
For human body generation, our method excels in both facial and hand synthesis. Our Face-Cos metric significantly surpasses that of the baseline model MimicMotion and is comparable to UniAnimate-DiT. This indicates that introducing object concepts did not compromise human generation quality.
Furthermore, VBench evaluations affirm that our generated videos exhibit superior subject consistency, background stability, and motion smoothness. These findings collectively validate the high visual quality and coherence of our approach.
5.2.2 Qualitative Results
Our qualitative results, illustrated in Fig. 6, demonstrate the efficacy of our approach. In these figures, red boxes denote insufficient object appearance fidelity, orange boxes highlight inaccuracies in hand synthesis, and green boxes indicate unrealistic anchor placements. The examples on the left pertain to stationary object interactions, whereas those on the right involve object displacement. Our method achieves superior preservation of both human and object appearances while faithfully capturing their dynamic behaviors.
MimicMotion and StableAnimator generate high-quality, pose-driven human images; however, they exhibit coupling artifacts, treating objects as static extensions of clothing rather than independent entities. When displacement occurs, objects either remain fixed in place or exhibit visual artifacts. UniAnimate-DiT maintains human appearance relatively well but suffers from object disappearance in certain frames and produces the least accurate hand synthesis. VACE struggles to preserve human appearance, leading to severe object deformation. Due to the complexity of hand-object interactions, all models except ours fail in generating realistic hand representations. For FlexiAct, we conducted 4,000 rounds of fine-tuning on the corresponding test videos, enabling it to fully learn object motion and appearance. While it successfully achieves object displacement, the overall visual quality remains suboptimal, and it fails to transfer interactive motions to novel subjects.
By leveraging HOI-appearance perception and HOI-motion injection, our method effectively preserves both the appearance and motion dynamics of humans and objects, producing high-fidelity human-object interaction videos.
@@ 5.3 Ablation Study
Figure 7: Ablation studies. Our modules improve the preservation of the object and its interactions with the hands. The error section has been enlarged in the lower right corner.
Validation of human-object dual adapter. We conducted detailed ablation experiments on human-object dual adapter. As shown in Table I. The Object-IoU metric remains superior to other models, which can be attributed to the detailed object trajectory guidance. However, the Object-CLIP score decreased, suggesting that merely injecting the latent representation of objects is insufficient for object perception, and the coupling with human subjects leads to a decline in face-cos. Fig. 7 presents qualitative results, demonstrating that the human and background significantly influence objects. The human-object dual adapter effectively learns object representations and decouples human and object features.
Validation of multi-view self-attention. As shown in Table I, the removal of the multi-view self-attention branch results in a slight decrease in object CLIP performance, as multi-view self-attention facilitates the extraction of more object information. Since the changes occur at a fine-grained level, they are difficult for CLIP to capture, leading to only a minimal drop in the score. Additionally, as illustrated in Fig. 7, when the object’s perspective shifts significantly, its logo undergoes deformation. Multi-view self-attention facilitates the extraction of more object information, effectively adapting to different viewpoints and accurately reconstructing the object’s appearance.
Validation of HOI-motion injection. We removed the hand 3D mesh information injection from the HOI-motion injection. As shown in Table I, the absence of detailed hand guidance resulted in a decrease in LMD (hand). Since errors are typically localized to specific fingers and are difficult for OpenPose to fully detect, the overall decline is not significant. The qualitative results illustrated in Fig. 7 show that the complexity of hand-object interactions, coupled with inaccurate hand pose guidance provided by DWPose, leads to the generation of artifacts. These artifacts become particularly noticeable when the hand is occluded by objects.
Validation of HOI-region reweighting loss. In conjunction with the HOI-region reweighting loss, we assigned higher weights to the hand-object regions during training to enhance the learning of objects and hands. As shown in Table I, insufficient emphasis on the hand-object regions causes the model to overfit to the human figures in the training set while failing to adequately learn object features, leading to inaccuracies in the appearance of both the object and the human figure. Fig. 7 further demonstrates that, without adequate region enhancement, the model struggles to capture fine object details effectively.
Figure 8: Ablation study of fine-tuning. After fine-tuning, the model can learn the texture details of objects.
Validation of Fine-Tuning During the fine-tuning phase, we focus on learning the intricate texture details of objects. Table I demonstrates that fine-tuning significantly enhances the Object-CLIP score. As illustrated in Fig. 8, when presented with a new object, the model successfully generates appropriate interactive actions, along with its contour and primary color. However, it lacks detailed texture representation. This observation indicates that our foundational model already possesses the capability for human-object interaction. Through fine-tuning, the model further develops the ability to perceive and represent fine-grained texture details.